ESX Quantum QE 900.4 continuously draw current

You could try changing the transistors in channel 1.

Is the idle current for the other channels constant whether they're hot or cold?

Is the falling side of the PS waveform free if the glitch it had previously?

The transformers in this amp are small. I don't know if going to 25k is going to be a problem. You'll have to drive the amp hard to see if there appears to be anything odd happening when hitting the hard notes.
 
I will replace the transistors, i have to get new ones first.

Channel 4 rises slightly, but no comparison to channel 1.

The glitch is gone, i think...I'll check it again tomorrow.

The transformers are getting hotter than before. I measured 60 degrees celsius, but at 4a idle.

You'll have to drive the amp hard

after each repair 🙂 not quite like in your tutorial but very similar
 
With the PS, I wanted to know if the frequency was normal and Ultra posted that he had seen it. OK.

There were a couple of things that concerned me and one was the glitch in the falling side of the drive. I don't know if the supply is regulated or not. It didn't appear to be in regulation from the waveforms you posted.

Were the PS FETs heating up at idle before you replaced the capacitor to lower the frequency?
 
OK. You may go back to the previous oscillation frequency, especially if the transformers are heating up.

Running the 3205s at twice the expected frequency may need a reduction in the gate resistor value but if they were running cool at idle and the supply isn't regulated, it wouldn't be absolutely necessary.

What B+ supply voltage are you getting at the B+ and ground terminals? The drain amplitude seemed a bit high.

If you have the full tutorial and didn't download a recent version, email me.
 
You may go back to the previous oscillation frequency, especially if the transformers are heating up.

If the spare parts are there i'll try ~35k, otherwise i'll go back up to 50k.

What B+ supply voltage are you getting at the B+ and ground terminals?

As set on the laboratory power supply unit.


The picture shows the Gate now. Glitch is disappeared, but another phenomenon was added.
 

Attachments

  • IMG_9755.JPEG
    IMG_9755.JPEG
    711.3 KB · Views: 45